# Quantum Hall Edges Beyond the Plasma Analogy

### v1

@@ Abstract
We demonstrate that the plasma analogy is generally unreliable at predicting the edge properties of quantum Hall (QH) states, as it fails to account for the local edge velocity. This discrepancy arises from a fundamental difference between QH droplets and Coulomb gases (CGs): the former are incompressible liquids subject to area-preserving deformations, while the latter are governed by electrostatics and thus involve conformal transformations. We illustrate this QH/CG mismatch across various examples and show its impact on physical quantities, measurable in both solid-state samples and quantum simulators. Specifically, we discuss fluctuations of local observables and their connection to state-of-the-art microwave absorption experiments.
@@ I Introduction
This work critically contrasts two commonly related physical systems. The first is a topological phase of matter famously responsible for the quantum Hall (QH) effect [1, 2, 3, 4, 5]: a two-dimensional (2D) incompressible quantum liquid of electrons in a strong perpendicular magnetic field. The second is a 2D Coulomb gas (CG) [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]: a classical statistical system of planar point charges interacting via a 2D electrostatic potential. Both systems are confined into droplets of finite area by trapping potentials, inherently present in any experiment.
@@
Our investigation is motivated by what is known as the plasma analogy, stating that QH systems are equivalent to suitable CGs. This correspondence underlies Laughlin’s theory of the fractional QH effect [5] and provides a handle on microscopic aspects that are otherwise out of reach due to strong electronic correlations. The analogy naturally also holds for integer QH droplets of free electrons in harmonic potentials, whose one-body wave functions coincide with those of a random matrix ensemble—the (elliptic) Ginibre ensemble [16, 17]. This includes the simplest, familiar case of isotropic QH droplets, whose wave functions have definite angular momentum. However, the plasma analogy remains untested for arbitrary potentials. This is a significant gap in the literature, given that realistic QH potentials are always disordered and anisotropic [18, 19, 20]. Besides, even in quantum simulators, where the potential can be controlled [21, 22, 23, 24, 25, 26, 27], it is desirable to know how much of the plasma analogy may be harnessed to model QH physics.
@@
The present paper puts the plasma analogy to the test. This relies on a recent semiclassical treatment of integer QH droplets in general anisotropic potentials [28], allowing for direct comparisons with CG predictions. The question is: Does the plasma analogy work for generic QH droplets? Strikingly, while reproducing bulk properties, the answer for edge phenomena is negative: QH droplets and CGs of identical shape almost always have different electronic correlations along the edge, resulting e.g. in unequal values for fluctuations of local observables. Such fluctuations can in principle be measured, at least indirectly, through microwave absorption experiments [29, 30, 31, 32, 33], so the implications of the QH/CG mismatch are both theoretical and experimental.
@@
Figure 1: QH droplets and CGs, shown in yellow along with a few level curves of the confining potential in the QH case. Both involve maps that send them on a disk $\mathbb{D}$, with different geometric constraints: symplectic (area-preserving) in QH, conformal (angle-preserving) in CG. Given the same initial droplet $\Omega$, such maps do not generally coincide, leading to different angle coordinates governing the emergent long-range edge correlations: the coordinate is $\theta$ for QH, $\alpha$ for CG, and $\varphi$ in the isotropic case.
@@
These differences between QH droplets and CGs should not come as a surprise. Since CGs ultimately rely on electrostatics, their physics is entirely dictated by their shape—albeit in a non-local manner. By contrast, QH physics is that of an incompressible fluid, so it depends not only on a droplet’s shape, but also on the local slope of the potential at the edge. This is the core of our argument. Abstractly, it can be phrased in terms of maps that send anisotropic droplets on their isotropic cousins, as in Fig. 1: the map is symplectic (area-preserving) for QH, but conformal (angle-preserving) for CG. It is still true that QH and CG predictions sometimes coincide, but this only occurs for fine-tuned setups, such as harmonic potentials [9]. Needless to say, no such fine-tuning is expected in solid-state experiments, where even just determining the confining potential is a challenge.
@@
The paper is organized as follows. Sec. II briefly reviews the links between QH physics and symplectic geometry on one hand, and CGs and conformal geometry on the other. This highlights the QH/CG discrepancy, whose implications are then examined in the remaining sections. Indeed, Sec. III illustrates the difference between QH and CG edge correlations through various explicit examples, and Sec. IV focuses on the consequences for fluctuations of local observables, some of which are evaluated using Monte-Carlo numerics on the CG side. Finally, we conclude in Sec. V. Further details on the mismatch, including brief reviews of QH droplets and CGs, are deferred to the Supplemental Material (SM) [34].
@@ II Quantum Hall is symplectic, Coulomb gas is conformal
Here we review the basics of QH and CG physics, and outline their respective links with symplectic and conformal geometries.
@@
This symplectic structure has important implications for QH systems. Namely, the guiding-center velocity (1) is tangent to equipotentials of $V$, and its norm
@@
is generally non-constant. One can nevertheless make the motion look isotropic in terms of well-chosen coordinates, at least when the equipotential has the topology of a circle. Indeed, let the equipotential have length $L$ and label its points by the arc length $s$ to define an angle
@@
where $s=0$ is arbitrary and $v(s)$ is the norm (2) at $s$. This is the canonical angle coordinate along the equipotential, unique up to rotation. It is conjugate to a canonical action coordinate $\mathcal{A}$, with dimensions of area, such that $\mathrm{d}\mathcal{A}\wedge\mathrm{d}\theta=2\pi\,\mathrm{d}x\wedge\mathrm{d}y$  [^1] In Hamiltonian mechanics, $\mathcal{A}$ measures the phase-space area enclosed by an energy level. Here, phase space is actually the non-commutative space of guiding-center positions, so “action” and “area” are synonymous.. Each value of $\mathcal{A}$ labels a level curve of $V$, so the potential can be written as $V(\textbf{x})=V_{0}(\mathcal{A}(\textbf{x}))$ for a strictly increasing function $V_{0}(\cdot)$ of $\mathcal{A}$ only, at least in a neighbourhood of the equipotential  [^2] Our normalization of the action coordinate $\mathcal{A}$ differs from that of the action coordinate $K$ in [28]. The two are related by $\mathcal{A}=2\pi\ell^{2}K$. Relatedly, $V_{0}(\cdot)$ here differs from that in [28] by a factor $2\pi$ in the argument.. These coordinates make the guiding center motion (1) uniform in the sense that
@@
When $V$ has a unique minimum surrounded by nested level curves, action-angle coordinates are globally well-defined on the plane and $\mathcal{A}$ can be chosen to be the actual area enclosed by an equipotential. For example, if $V$ is isotropic, then $\mathcal{A}=\pi r^{2}$ and $\theta=\varphi$ in polar coordinates.
@@
that converts the anisotropic potential $V$ into its isotropic cousin $V_{0}(\mathcal{A})$, as in Fig. 1. Here we write the map as if it was globally well-defined, which is generally neither true nor required: all one needs is that it sends a neighbourhood of an equipotential on a neighbourhood of a circle in a smooth, bijective, and area-preserving manner. In the QH case, the equipotential of interest will be the edge of a droplet, whose topology is always that of a circle no matter how disordered the bulk potential is.
These classical structures carry over to quantum physics, including many-body states at low energy. Position operators projected to the lowest Landau level (LLL) fail to commute, and the position of a coherent state in the LLL satisfies Eq. (1) in the semiclassical limit of weak potentials and strong magnetic fields [37, 38, 39]. Relatedly, semiclassical eigenstates in the LLL are localized on equipotentials, each enclosing a quantized area $\mathcal{A}=2\pi m\ell^{2}$ where $\ell\equiv\sqrt{\hbar/|qB|}$ is the magnetic length and $m$ is a (large) positive integer [40, 41, 28]. Extrapolating to a many-body system of $N\gg 1$ free electrons in the LLL is straightforward: they occupy the domain $\Omega$ where $V(\textbf{x})\leq E_{\textrm{F}}$ for some Fermi energy $E_{\textrm{F}}$, forming a quantum Hall (QH) droplet with area $\mathcal{A}=2\pi N\ell^{2}$ and uniform bulk density $\rho=1/2\pi\ell^{2}$. A similar picture is expected for fractional QH droplets confined by a weak potential [2, 5], then with total area $\mathcal{A}=2\pi N\ell^{2}/\nu$ and bulk density $\rho=\nu/2\pi\ell^{2}$, both depending on the filling fraction $\nu$.
@@
$\frac{W(\textbf{x})}{2}-\int_{\Omega}\mathrm{d}\textbf{y}\,\rho(\textbf{y})\log|\textbf{x}-\textbf{y}|=\text{const}$ (8)
valid for all $\textbf{x}\in\Omega$, where $\mathrm{d}\textbf{y}\equiv\mathrm{d}y^{1}\mathrm{d}y^{2}$. Taking the Laplacian of both sides gives $\rho=\nabla^{2}W/4\pi$ inside the droplet and $\rho=0$ outside, which fixes the density. In turn, plugging $\rho=\nabla^{2}W/4\pi$ back into Eq. (8) implicitly determines $\Omega$ through a (generally complicated) inverse problem. The droplet’s shape is thus determined by the potential $W$ but, in contrast to QH droplets, it is also strongly affected by Coulomb interactions. (This is why our use of $V$ versus $W$ is important to distinguish QH and CG potentials.) A much easier task is to first choose $\Omega$ and a normalized $\rho$ therein, then reverse-engineer the potential $W$ in $\Omega$ through Eq. (8). In fact, this is precisely how CGs with uniform bulk density are meant to be used in the plasma analogy.
@@
A simple way to understand the relevance of conformal geometry outside CGs and on their edge is to study correlations. In that context, the fundamental object is the (connected) two-point function of the field (9), which can be determined from linear response. Indeed, suppose a (small) charge $q$ is inserted at some position x, so that the energy $E$ in Eq. (7) is modified to $E+q\Phi(\textbf{x})$. Denoting by $\braket{\cdot}^{\textbf{x}}$ the corresponding expectation value, linear response yields
@@
maps the exterior of the droplet on the exterior of the disk  [^4] For comparison with QH droplets, we assume that the disk $\mathbb{D}$ and the droplet $\Omega$ in (11) have the same area. This entails no loss of generality, since it can be ensured by composing $G(\textbf{x})$ with an overall dilation that does not affect the angle $\alpha$ introduced below the map (11)., whereupon the Green’s function outside $\Omega$ is $K(\textbf{x},\textbf{y})=K_{0}(G(\textbf{x}),G(\textbf{y}))$. This provides a natural “conformal” angle coordinate, which we denote $\alpha$: for $\textbf{x}\in\partial\Omega$, it is the polar angle of the vector $G(\textbf{x})$.
We stress that conformal maps such as (11) are extremely rigid, since they have to be invertible and holomorphic (in the coordinate $x^{1}+\mathrm{i}x^{2}$) everywhere outside $\Omega$. Indeed, provided the map (11) sends infinity on itself, the Riemann mapping theorem ensures that it is unique up to an overall rotation. Put differently, all correlations in a CG are fully determined by the shape of the droplet $\Omega$.
Figure 2: Monte-Carlo simulation of the norm of the total electric field $\mathbf{E}(\textbf{y})=-\nabla_{\textbf{y}}\bigl{(}\braket{\Phi(\textbf{y})}^{\textbf{x}}-\braket{\Phi(\textbf{y})}-\log|\textbf{x}-\textbf{y}|\bigr{)}$ for $N=256$ particles and potential $W(x,y)=x^{2}+y^{2}-(x^{3}-3xy^{2})/10$, shown as a function of $\textbf{y}=(y^{1},y^{2})$. (The potential is polynomial so as to give an anisotropic CG at minimal numerical cost, but is otherwise unimportant.) Left: Fixed unit charge at $\textbf{x}\in\Omega$; its effect is screened over short distances. Right: Fixed unit charge at $\textbf{x}\notin\Omega$, in which case the electric potential is a long ranged Green’s function. In both plots, the electric field vanishes inside $\Omega$ and is discontinuous on $\partial\Omega$, signaling charge accumulation on the boundary.
An important application of the conformal map (11) is provided by density-density correlations, especially since they can be compared to their counterparts in QH droplets. While bulk correlations are short ranged due to screening [6], edge correlations are long ranged. They can be predicted from electrostatics: they are, in fact, correlations of the edge charge density produced at the boundary of the CG by the discontinuous electric field; see again Fig. 2. The resulting correlator behaves as
@@
The mismatch. The difference between QH droplets and CGs can be summarized in terms of the maps (5) and (11). Consider a (simply connected) domain $\Omega$ in the plane and compare two scenarios: one where it is occupied by a QH droplet in a potential $V$, and another where it is occupied by a CG in a potential $W$. Both systems share the same boundary $\partial\Omega$. The density correlator between points $\textbf{x}_{1}$ and $\textbf{x}_{2}$ on $\partial\Omega$ is given by Eq. (6) with $\theta_{i}=\theta(\textbf{x}_{i})$ for QH, and by Eq. (12) with $\alpha_{i}=\alpha(\textbf{x}_{i})$ for CG. The coordinates $\theta_{i}$ and $\alpha_{i}$ are respectively derived from the area-preserving map (5) and the angle-preserving map (11). These maps usually differ, except in a few highly specific cases, causing the discrepancy between Eqs. (6) and (12) despite their superficial similarity. Physically, this is because electrostatics is only sensitive to the shape of a CG, while QH physics is sensitive to both the shape and the local gradient of the confining potential at the edge.
This mismatch is our main statement, so let us rephrase it for later reference. Namely, the maps (5) and (11) yield two parametrizations of the edge:
@@
These typically differ, in the sense that the 1D diffeomorphism $\alpha\circ\theta^{-1}:S^{1}\to S^{1}$ is not just a rotation $\varphi\mapsto\varphi+\textrm{const}$ (except in rare special cases). In the remainder of this work, we verify this claim in several situations of interest and point out its experimental ramifications.

### v2

@@ Abstract
We demonstrate that the widely used plasma analogy is unreliable at predicting edge properties of quantum Hall states. This discrepancy arises from a fundamental difference between quantum Hall droplets and plasmas (Coulomb gases): the former are incompressible liquids subject to area-preserving deformations, while the latter are governed by electrostatics and thus involve conformal transformations. Consequently, the plasma analogy fails at the edge, except in fine-tuned geometries, as it does not account for the emergent local edge velocity. We quantitatively show how the analogy’s failure affects physical quantities, such as fluctuations of local observables and absorption rates in microwave spectroscopy, measurable in both solid-state experiments and quantum simulators.
@@ I Introduction
This paper critically contrasts two commonly related physical systems. The first is a topological phase of matter famously responsible for the quantum Hall (QH) effect [1, 2, 3, 4, 5]: a two-dimensional (2D) incompressible quantum liquid of electrons in a strong perpendicular magnetic field. The second is a 2D Coulomb gas (CG) [6, 7, 8, 9, 10, 11, 12, 13, 14, 15]: a classical statistical system of planar point charges interacting via a 2D electrostatic potential. Both systems are confined into droplets of finite area by trapping potentials, inherently present in any experiment.
@@
Our investigation is motivated by what is known as the plasma analogy, stating that QH systems are equivalent to suitable CGs. This correspondence underlies Laughlin’s theory of the fractional QH effect [5] and provides a handle on microscopic aspects that are otherwise out of reach due to strong electronic correlations. The analogy naturally also holds for integer QH droplets of free electrons in harmonic potentials, whose one-body wave functions coincide with those of a random matrix ensemble—the (elliptic) Ginibre ensemble [16, 17]. This includes the simplest, familiar case of isotropic QH droplets, whose wave functions have definite angular momentum. However, the plasma analogy remains untested for arbitrary potentials. This is a significant gap in the literature, given that realistic QH potentials are always disordered and anisotropic [18, 19, 20]. Additionally, even in quantum simulators, where the potential can be controlled [21, 22, 23, 24, 25, 26, 27], it is desirable to know how much of the plasma analogy may be harnessed to model QH physics.
@@
In the present paper, we put the plasma analogy to the test. This relies on a recent semiclassical treatment of integer QH droplets in general anisotropic potentials [28], allowing for direct comparisons with CG predictions. The question is: Does the plasma analogy work for generic QH droplets? Strikingly, while reproducing bulk properties, the answer for edge phenomena is negative: QH droplets and CGs of identical shape almost always have different electronic correlations along the edge, resulting e.g. in unequal values for fluctuations of local observables. Such fluctuations can be measured, at least indirectly, through microwave absorption experiments [29, 30, 31, 32, 33], so the implications of the QH/CG mismatch are both theoretical and experimental.
@@
Figure 1: QH droplets and CGs, shown in yellow along with a few level curves of the confining potential in the QH case. Both involve maps that send them on a disk $\mathbb{D}$, but with different geometric constraints, namely symplectic (area-preserving) in QH and conformal (angle-preserving) in CG. Given the same initial droplet $\Omega$, such maps do not generally coincide, leading to different angle coordinates governing the emergent long-range edge correlations: the coordinate is $\theta$ for QH, $\alpha$ for CG, and $\varphi$ in the isotropic case.
@@
These differences between QH droplets and CGs should not come as a surprise. Since CGs ultimately rely on electrostatics, their physics is entirely dictated by their shape—albeit in a nonlocal manner. By contrast, QH physics is that of an incompressible fluid, so it depends not only on a droplet’s shape but also on the local slope of the potential at the edge. This is the core of our argument. Abstractly, it can be phrased in terms of maps that send anisotropic droplets on their isotropic cousins, as in Fig. 1: the map is symplectic (area-preserving) for QH, but conformal (angle-preserving) for CG. It is still true that QH and CG predictions sometimes coincide, but this only occurs for fine-tuned setups, such as harmonic potentials [9]. Needless to say, no such fine-tuning is expected in solid-state experiments, where even just determining the confining potential is a challenge.
@@
The paper is organized as follows. Sec. II briefly reviews the links between QH physics and symplectic geometry on one hand, and CGs and conformal geometry on the other. This highlights the QH/CG discrepancy, whose implications are then examined in the remaining sections. Indeed, Sec. III illustrates the difference between QH and CG edge correlations through various explicit examples, and Sec. IV focuses on the consequences for fluctuations of local observables, some of which are evaluated using Monte Carlo numerics on the CG side. Finally, we conclude in Sec. V. Further details on the mismatch, including brief reviews of QH droplets and CGs, are deferred to the Supplemental Material (SM) [34].
@@ II Quantum Hall is symplectic, Coulomb gas is conformal
Here, we review the basics of QH and CG physics and outline their respective links with symplectic and conformal geometries.
@@
This symplectic structure has important implications. Namely, the guiding-center velocity (1) is tangent to equipotentials of $V$, and its norm
@@
is generally nonconstant. One can nevertheless make the motion look isotropic in terms of well-chosen coordinates, at least when the equipotential has the topology of a circle. Indeed, let the equipotential have length $L$ and label its points by the arc length $s$ to define an angle
@@
where the lower integration limit is arbitrary and $v(s)$ is the norm (2) at $s$. This is the canonical angle coordinate along the equipotential, unique up to rotation. It is conjugate to a canonical action coordinate $\mathcal{A}$, with dimensions of area, such that $\mathrm{d}\mathcal{A}\wedge\mathrm{d}\theta=2\pi\,\mathrm{d}x\wedge\mathrm{d}y$  [^1] In Hamiltonian mechanics, $\mathcal{A}$ measures the phase-space area enclosed by an energy level. Here, phase space is the noncommutative space of guiding-center positions, so action and area are synonymous.. Each value of $\mathcal{A}$ labels a level curve of $V$, so the potential can be written as $V(\textbf{x})=V_{0}(\mathcal{A}(\textbf{x}))$ for a strictly increasing function $V_{0}(\cdot)$ of $\mathcal{A}$ only, at least in a neighborhood of the equipotential  [^2] Our normalization of the action coordinate $\mathcal{A}$ differs from that of the action coordinate $K$ in [28]. The two are related by $\mathcal{A}=2\pi\ell^{2}K$. Relatedly, $V_{0}(\cdot)$ here differs from that in [28] by a factor $2\pi$ in the argument.. These coordinates make the guiding center motion in Eq. (1) uniform in the sense that
@@
When $V$ has a unique minimum surrounded by nested level curves, action-angle coordinates are globally well defined on the plane, and $\mathcal{A}$ can be chosen to be the actual area enclosed by an equipotential. For example, if $V$ is isotropic, then $\mathcal{A}=\pi r^{2}$ and $\theta=\varphi$ in polar coordinates.
@@
that converts the anisotropic potential $V$ into its isotropic cousin $V_{0}(\mathcal{A})$, as in Fig. 1. Here, we write the map as if it was globally well defined, which is generally neither true nor required: all one needs is that it sends a neighborhood of an equipotential on a neighborhood of a circle in a smooth, bijective, and area-preserving manner. In our case, the equipotential of interest will be the edge of a droplet, whose topology is always that of a circle no matter how disordered the bulk potential is.
Turning to quantum physics, the corresponding problem of an electron in a strong magnetic field with a harmonic potential dates back to [37, 38] and is solvable using special functions. More generally, for any confining potential $V$, the above classical structures carry over to the quantum case, including many-body states at low energy. Position operators projected to the lowest Landau level (LLL) fail to commute, and the position of a coherent state in the LLL satisfies Eq. (1) in the semiclassical limit of weak potentials and strong magnetic fields [39, 40, 41]. Relatedly, semiclassical eigenstates in the LLL are localized on equipotentials, each enclosing a quantized area $\mathcal{A}=2\pi m\ell^{2}$, where $\ell\equiv\sqrt{\hbar/|qB|}$ is the magnetic length and $m$ is a (large) positive integer [42, 43, 28]. Extrapolating to a many-body system of $N\gg 1$ free electrons in the LLL is straightforward: they occupy the domain $\Omega$ where $V(\textbf{x})\leq E_{\textrm{F}}$ for some Fermi energy $E_{\textrm{F}}$, forming a quantum Hall (QH) droplet with area $\mathcal{A}=2\pi N\ell^{2}$ and uniform bulk density $\rho=1/2\pi\ell^{2}$. A similar picture is expected for fractional QH droplets confined by a weak potential [2, 5], then with total area $\mathcal{A}=2\pi N\ell^{2}/\nu$ and bulk density $\rho=\nu/2\pi\ell^{2}$, both depending on the filling fraction $\nu$.
@@
$\frac{W(\textbf{x})}{2}-\int_{\Omega}\mathrm{d}\textbf{y}\,\rho(\textbf{y})\log|\textbf{x}-\textbf{y}|=\text{const},$ (8)
valid for all $\textbf{x}\in\Omega$, where $\mathrm{d}\textbf{y}\equiv\mathrm{d}y^{1}\mathrm{d}y^{2}$. Taking the Laplacian of both sides gives $\rho=\nabla^{2}W/4\pi$ inside the droplet and $\rho=0$ outside, which fixes the density. In turn, plugging $\rho=\nabla^{2}W/4\pi$ back into Eq. (8) implicitly determines $\Omega$ through a (generally complicated) inverse problem. The droplet’s shape is thus determined by the potential $W$, but in contrast to QH droplets, it is also strongly affected by Coulomb interactions. (This is why our use of $V$ vs $W$ is important to distinguish QH and CG potentials.) A much easier task is to first choose $\Omega$ and a normalized $\rho$ therein, then reverse-engineer the potential $W$ in $\Omega$ through Eq. (8). In fact, this is precisely how CGs with uniform bulk density are meant to be used in the plasma analogy.
@@
A simple way to understand the relevance of conformal geometry outside CGs and on their edge is to study correlations. In that context, the fundamental object is the (connected) two-point function of the field in Eq. (9), which can be determined from linear response. Indeed, suppose a (small) charge $q$ is inserted at some position x, so that the energy $E$ in Eq. (7) is modified to $E+q\Phi(\textbf{x})$. Denoting by $\braket{\cdot}^{\textbf{x}}$ the corresponding expectation value, linear response yields
@@
maps the exterior of the droplet on the exterior of the disk  [^4] For comparison with QH droplets, we assume that the disk $\mathbb{D}$ and the droplet $\Omega$ in Eq. (11) have the same area. This entails no loss of generality, since it can be ensured by composing $G(\textbf{x})$ with an overall dilation that does not affect the angle $\alpha$ introduced below the map (11)., whereupon the Green’s function outside $\Omega$ is $K(\textbf{x},\textbf{y})=K_{0}(G(\textbf{x}),G(\textbf{y}))$. This provides a natural conformal angle coordinate, which we denote $\alpha$: for $\textbf{x}\in\partial\Omega$, it is the polar angle of the vector $G(\textbf{x})$.
We stress that conformal maps such as in Eq. (11) are extremely rigid since they must be invertible and holomorphic (in the coordinate $x^{1}+\mathrm{i}x^{2}$) everywhere outside $\Omega$. Indeed, provided the map (11) sends infinity on itself, the Riemann mapping theorem ensures that it is unique up to an overall rotation. Put differently, all correlations in a CG are fully determined by the shape of the droplet $\Omega$.
Figure 2: Monte Carlo simulation of the norm of the total electric field $\mathbf{E}(\textbf{y})=-\nabla_{\textbf{y}}\bigl{(}\braket{\Phi(\textbf{y})}^{\textbf{x}}-\braket{\Phi(\textbf{y})}-\log|\textbf{x}-\textbf{y}|\bigr{)}$ for $N=256$ particles and potential $W(x,y)=x^{2}+y^{2}-(x^{3}-3xy^{2})/10$, shown as a function of $\textbf{y}=(y^{1},y^{2})$. (The potential is polynomial so as to give an anisotropic CG at minimal numerical cost, but it is otherwise unimportant.) Left: Fixed unit charge at $\textbf{x}\in\Omega$; its effect is screened over short distances. Right: Fixed unit charge at $\textbf{x}\notin\Omega$, in which case the electric potential is a long-range Green’s function. In both plots, the electric field vanishes inside $\Omega$ and is discontinuous on $\partial\Omega$, signaling charge accumulation on the boundary.
An important application of the conformal map (11) is provided by density-density correlations, especially since they can be compared with their counterparts in QH droplets. While bulk correlations are short-ranged due to screening [6], edge correlations are long-ranged. They can be predicted from electrostatics: they are, in fact, correlations of the edge charge density produced at the boundary of the CG by the discontinuous electric field; see again Fig. 2. The resulting correlator behaves as
@@
Mismatch. The difference between QH droplets and CGs can be summarized in terms of the maps in Eqs. (5) and (11). Consider a (simply connected) domain $\Omega$ in the plane, and compare two scenarios: one where it is occupied by a QH droplet in a potential $V$, and another where it is occupied by a CG in a potential $W$. Both systems share the same boundary $\partial\Omega$. The density correlator between points $\textbf{x}_{1}$ and $\textbf{x}_{2}$ on $\partial\Omega$ is given by Eq. (6) with $\theta_{i}=\theta(\textbf{x}_{i})$ for QH, and by Eq. (12) with $\alpha_{i}=\alpha(\textbf{x}_{i})$ for CG. The coordinates $\theta_{i}$ and $\alpha_{i}$ are respectively derived from the area-preserving map (5) and the angle-preserving map (11). These maps usually differ, except in a few highly specific cases, causing the discrepancy between Eqs. (6) and (12) despite their superficial similarity. Physically, this is because electrostatics is only sensitive to the shape of a CG, while QH physics is sensitive to both the shape and the local gradient of the confining potential at the edge.
This mismatch is our main statement, so let us rephrase it for later reference. Namely, the maps in Eqs. (5) and (11) yield two parameterizations of the edge:
@@
These typically differ, in the sense that the one-dimensional diffeomorphism $\alpha\circ\theta^{-1}:S^{1}\to S^{1}$ is not just a rotation $\varphi\mapsto\varphi+\textrm{const}$ (except in rare special cases). In the remainder of this paper, we verify this claim in several situations of interest and point out its experimental ramifications. We stress that this does not affect the universal correspondence between bulk topology and edge modes [50]: indeed, as shown in [28], the emergent effective edge theory is unaffected by our deformations of the sample’s geometry (at least to leading order in $N$).